OpenStudy (anonymous):

Find the slope between points (1, 8) and (9, 1) i need the forumla

OpenStudy (nurali):

m=y2-y1/x2-x1

OpenStudy (anonymous):

ok... rise over run

OpenStudy (nurali):

(1, 8) and (9, 1)=(x1,y1) and (x2,y2)

OpenStudy (anonymous):

\[\frac{ 8 }{ -7 }\]

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

i mean

OpenStudy (anonymous):

8 = rise over negative 7 because when your on 8 and you go ⬇ it has to be a negative before the number so \[\frac{ 8 }{ -7 }\]

OpenStudy (nurali):

\[m=\frac{ 1-8 }{ 9-1 }\]

OpenStudy (anonymous):

i did it backwards huh?

OpenStudy (anonymous):

yup i did ... wasn't paying attention